Solve the following equations and check your results  2x/3 + 1 = 7x/15 + 3 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2x/3+1=7x/15+3
⇒2x/3–7x/15=3–1
⇒(10x–7x)/15=2
⇒3x=2×15
⇒3x=30
⇒x=30/3
⇒x=10
Putting the value of x in RHS and LHS we get,

Thanks